# Settling Constant Regrets in Linear Markov Decision Processes

### v1

@@ 6.1 Technical challenges
6.2 A novel approach: Cert-LinUCB
We introduce the technical details in designing our new subroutine, Cert-LinUCB, to tackle Challenge 1. In the addition of using ‘local quantification’ which ensure the quantification error of each phase $l$ depend on the local phase $\widetilde{\mathcal{O}}(l)$ instead of the global parameter $\log K$, Algorithm 2 eliminates the $\log K$ dependence from $\log|\mathcal{V}|$ by employing certified estimator.
Considering the concentration term we need to control for each phase $l$:
$\Big{\|}\textstyle{\sum_{\tau\in\mathcal{C}_{h,l}^{k}}}\bm{\phi}^{\tau}_{h}\big{[}\widehat{V}_{h+1}^{k}(s^{\tau})-\mathbb{E}[\widehat{V}_{h+1}^{k}(s^{\tau})]\big{]}\Big{\|}_{(\mathbf{U}^{k}_{h,l})^{-1}},$
Lemma 6.1 (Lemma C.4, informal).
Let $\widehat{V}_{h+1,l_{+}}^{k}$ be the output of Algorithm 2 terminated at phase $L=l_{+}$, then with probability is at least $1-2\delta$,
To apply Lemma 6.1, it is essential to bound the distance between $\widehat{V}_{h+1,l_{+}}^{k}$ and $\widehat{V}_{h+1}^{k}$. For this purpose, we maintain a monotonic sequence of the optimistic value function $\widehat{V}_{h+1,l}^{k}$ and the pessimistic value function $\widecheck{V}_{h+1,l}^{k}$, ensuring that
$\widecheck{V}_{h+1,1}^{k}(s)\leq\widecheck{V}_{h+1,2}^{k}(s)\leq\cdots\leq\widecheck{V}_{h+1,l^{k}_{h}(s)-1}^{k}(s)$
$\leq\widehat{V}_{h+1}^{k}(s)=\widehat{V}_{h+1,l^{k}_{h}(s)-1}^{k}(s)\leq\cdots\leq\widehat{V}_{h+1,2}^{k}(s)\leq\widehat{V}_{h+1,1}^{k}(s).$ (6.2)
This monotonicity is guaranteed according to Line 11 in Cert-LinUCB, where the process is terminated once (6.2) is violated. As a result, we can control the distance between $\widehat{V}_{h+1,l_{+}}^{k}$ and $\widehat{V}_{h+1}^{k}$ as the following lemma.
Lemma 6.2 (Lemma C.2, informal).
There is a faithful extension of $\widehat{V}_{h+1,l_{+}}^{k}$ to every $l_{+}\in\mathbb{N}^{+}$ that $|\widehat{V}^{k}_{h}(s)-\widehat{V}^{k}_{h,l_{+}}(s)|\leq 6\cdot 2^{-l_{+}}$ always holds.
Following these results, combining Lemma 6.2 and Lemma 6.1 together obtains a local concentration bound for each phase $l$ that is independent of $\log K$ when choosing $l_{+}=l+\widetilde{\mathcal{O}}(\log(ld))$.
Lemma 6.3 (Lemma C.5, informal).
With probability at least $1-2\delta$, for any $(k,h,l)\in[K]\times[H]\times\mathbb{N}^{+}$:
$\Big{\|}\textstyle{\sum_{\tau\in\mathcal{C}^{k-1}_{h,l}}}\bm{\phi}^{\tau}_{h}\big{(}[\mathbb{P}_{h}\widehat{V}^{k}_{h+1}](s^{\tau}_{h},a^{\tau}_{h})-\widehat{V}^{k}_{h+1}(s^{\tau}_{h+1})\big{)}\Big{\|}_{(\mathbf{U}^{k}_{h,l})^{-1}}\leq 1.1\gamma_{l}.$
As a result of our improved concentration analysis, we can achieve a local estimation error for the estimated value function in each phase $l$:
Lemma 6.4 (Lemma C.6, informal).
With high probability, for any $(k,h,s)\in[K]\times[H]\times{\mathcal{S}},l\in[l^{k}_{h}(s)-f^{k}_{h}(s)],a_{l}\in\mathcal{A}^{k}_{h,l}(s)$, we have $\big{|}Q^{k}_{h,l}(s,a)-[\mathbb{B}_{h}\widehat{V}^{k}_{h+1}](s,a)\big{|}\leq 2\cdot 2^{-l}+\mathcal{O}(\sqrt{d}H\zeta).$
Lemma 6.4 bounds the estimation error for any state $s$ by the phase $l_{h}^{k}(s)$ where $l_{h}^{k}(s)$ indicates the layer at which Algorithm 2 terminates. As the early phases cannot provide sufficient accuracy due to a lack of data, we also need to analyze the conditions under which Line 11 is triggered in Algorithm 2.
Lemma 6.5 (Lemma C.8, informal).
Lemma 6.5 delivers a clear message: the trigger of Line 11 is related to the misspecification level $\zeta$. In the well-specified setting, Line 11 will never be triggered ($l\geq\infty$). When the misspecification level is large, then Line 11 will be more likely triggered, indicating it’s harder to get higher precise estimation via higher $l_{h}^{k}(s)$, according to Lemma 6.4.
In order to bound the number of suboptimality gap taken by Cert-LSVI-UCB, we start with the following standard decomposition
$V_{h}^{*}(s^{k}_{h})-$ $V^{\pi^{k}}_{h}(s^{k}_{h})=\big{(}V_{h}^{*}(s^{k}_{h})-\widehat{V}^{k}_{h}(s^{k}_{h})\big{)}+\big{(}\widehat{V}^{k}_{h}(s^{k}_{h})-V^{\pi^{k}}_{h}(s^{k}_{h})\big{)}$
$=\big{(}V_{h}^{*}(s^{k}_{h})-\widehat{V}^{k}_{h}(s^{k}_{h})\big{)}+\textstyle{\sum_{h^{\prime}=h}^{H}}\Big{(}\widehat{V}^{k}_{h^{\prime}}(s^{k}_{h^{\prime}})-[\mathbb{B}_{h}\widehat{V}^{k}_{h^{\prime}+1}]\big{(}s^{k}_{h^{\prime}},\pi^{k}_{h^{\prime}}(s^{k}_{h^{\prime}})\big{)}\Big{)}+\textstyle{\sum_{h^{\prime}=h}^{H}}\eta^{k}_{h^{\prime}}.$
where $\eta^{k}_{h}=[\mathbb{P}_{h}(\widehat{V}^{k}_{h+1}-V_{h+1}^{\pi^{k}})]\big{(}s^{k}_{h},\pi^{k}_{h}(s^{k}_{h})\big{)}-\big{(}\widehat{V}^{k}_{h+1}(s^{k}_{h+1})-V^{\pi^{k}}_{h+1}(s^{k}_{h+1})\big{)}$ is a zero-mean random variable induced by the transition kernel. We can bound each of the factors using standard regret analysis on the basis created by combining Lemma 6.4 and Lemma 6.5:
  1. 1.
    Global underestimation error $V_{h}^{*}(s^{k}_{h})-\widehat{V}^{k}_{h}(s^{k}_{h})$. (see Lemma D.11)
  2. 2.
    Local overestimation error $\widehat{V}^{k}_{h}(s^{k}_{h})-[\mathbb{B}_{h}\widehat{V}^{k}_{h+1}](s^{k}_{h^{\prime}},\pi^{k}_{h}\big{(}s^{k}_{h})\big{)}$. (see Lemma D.12)
  3. 3.
    Transition noise $\eta^{k}_{h}$. (see Lemma D.14)
In general, we can reach the following results, which provide an local regret upper bound for arbitrary index subsets which is independent from the number of total episodes.
Lemma 6.6 (Lemma D.15, Informal).
With high probability, for any index set $\mathcal{K}$ and any $\varepsilon$ that is comparably large against $\zeta$, it satisfies that
$\textstyle{\sum_{k\in\mathcal{K}}}\big{(}V_{h}^{*}(s^{k}_{h})-V^{\pi^{k}}_{h}(s^{k}_{h})\big{)}\leq 0.49|\mathcal{K}|\varepsilon+\widetilde{\mathcal{O}}\big{(}d^{3}H^{4}\varepsilon^{-1}+\sqrt{H^{3}|\mathcal{K}|}\big{)}.$
Note that for index set $\mathcal{K}=\big{[}k:V_{h}^{*}(s^{k}_{h})-V^{\pi^{k}}_{h}(s^{k}_{h})\geq\varepsilon\big{]}$, the regret enjoys a trivial lower bound that $\sum_{k\in\mathcal{K}}\big{(}V_{h}^{*}(s^{k}_{h})-V^{\pi^{k}}_{h}(s^{k}_{h})\big{)}\geq|\mathcal{K}|\varepsilon$. We thus can reach the following result.
Lemma 6.7 (Lemma C.12, Informal).
Remark 6.8.
He et al. (2021b) achieved a uniform-PAC bound for (well-specified) linear MDP, which states as
$w.h.p.,\forall\epsilon>0,\textstyle{\sum_{k=1}^{\infty}}\operatorname{\mathds{1}}[V_{1}^{*}(s_{1}^{k})-V_{1}^{\pi^{k}}(s_{1}^{k})\geq\epsilon]\leq\widetilde{\mathcal{O}}(d^{3}H^{5}\epsilon^{-2}),$ (6.3)
comparing (6.3) with Lemma 6.7 on well-specified setting where $\zeta=0$, one can find that our result is better than He et al. (2021b) under stronger condition: Lemma 6.7 ensures this uniform-PAC result under all stage $h\in[H]$ while He et al. (2021b) only ensure the initial statement. Due to the randomness of MDP discussed in Challenge 2, the guarantee for $h=1$ cannot be easily generated to any $h\in[H]$. Second, our result $\widetilde{\mathcal{O}}(d^{3}H^{4}\epsilon^{-2})$ improves He et al. (2021b) by a factor $H$. This is because a more efficient data selection strategy which we will discuss in detail in Appendix B.1.
6.3 Settling the gap between $V^{*}-V^{\pi}$ and $V^{*}-Q^{*}$
According to Challenge 2, the regret in episodes where $V_{1}^{*}(s^{k}_{1})-V^{\pi^{k}}_{1}(s^{k}_{1})\leq\Delta$ is non-zero since the minimal suboptimality gap assumption $\Delta$ only guarantees $\Delta^{k}_{h}=V^{*}_{h}(s^{k}_{h})-Q^{*}_{h}\big{(}s^{k}_{h},\pi^{k}_{h}(s^{k}_{h})\big{)}\notin(0,\Delta)$ but put no restrictions on $V^{*}_{h}(s^{k}_{h})-V^{\pi^{k}}_{h}(s^{k}_{h})$.
@@ 6.3 Settling the gap between $V^{*}-V^{\pi}$ and $V^{*}-Q^{*}$
Notice that the regret $V^{*}_{h}(s_{h})-V^{\pi^{k}}_{h}(s_{h})$ in episode $k$ is the expectation of cumulative suboptimality gap $\sum_{h=1}^{H}\Delta^{k}_{h}$ taking over trajectory $\{s^{k}_{h}\}_{h=1}^{H}$. In addition, the variance of the random variable can be self-bounded according to
@@
As Freedman inequality (Lemma H.5) implies $\sum_{t=1}^{T}\operatorname{{\rm Var}}[\eta^{t}]\leq aC$ and $\sum_{t=1}^{T}\operatorname{{\rm Var}}[\eta^{t}]\leq vC$ only happens with small probability for every $C$ with proper constant $a$ and $v$, together with a union bound statement $C$, we can reach the following statement indicates the cumulative regret can be upper bounded using the cumulative suboptimality gap:
@@
$\textstyle{\sum_{k=1}^{K}}\big{(}V^{*}_{h}(s_{h})-V^{\pi^{k}}_{h}(s_{h})\big{)}\leq\widetilde{\mathcal{O}}\Big{(}\textstyle{\sum_{k=1}^{K}}\textstyle{\sum_{h=1}^{H}}\Delta^{k}_{h}+H^{2}\Big{)}.$
In addition, since the loss on the state-value function $V^{*}_{h}(s^{k}_{h})-V^{\pi^{k}}_{h}(s^{k}_{h})$ is a upper bound for the sub-optiomality gap $\Delta^{k}_{h}$, we are able to show that the cumulative suboptimality gap is constantly bounded when the minimal suboptimality gap is sufficiently large as in Lemma C.13:
$\textstyle{\sum_{k=1}^{K}}$ $\textstyle{\sum_{h=1}^{H}\Delta^{k}_{h}=\sum_{k=1}^{K}\sum_{h=1}^{H}\Big{(}\Delta\cdot\operatorname{\mathds{1}}\Big{[}\Delta^{k}_{h}\geq\Delta\Big{]}+\int_{\Delta}^{H}\operatorname{\mathds{1}}\Big{[}\Delta^{k}_{h}\geq\varepsilon\Big{]}{\,\mathrm{d}}\varepsilon\Big{)}}$
$\leq\textstyle{\sum_{k=1}^{K}\sum_{h=1}^{H}\Big{(}\Delta\cdot\operatorname{\mathds{1}}\Big{[}V^{*}_{h}(s^{k}_{h})-V^{\pi^{k}}_{h}(s^{k}_{h})\geq\Delta\Big{]}+\int_{\Delta}^{H}\operatorname{\mathds{1}}\Big{[}V^{*}_{h}(s^{k}_{h})-V^{\pi^{k}}_{h}(s^{k}_{h})\geq\varepsilon\Big{]}{\,\mathrm{d}}\varepsilon\Big{)}}.$
Note that the final summation can be upper bounded by $\widetilde{\mathcal{O}}\big{(}d^{3}H^{5}\Delta^{-2})$ using Lemma 6.7. Together with Lemma 6.9, we reach the desired statement that Cert-LSVI-UCB achieves constant regret bound when the misspecification is sufficiently small against the minimal suboptimality gap.
7 Conclusion
In this work, we proposed a new algorithm, called certified estimator, for reinforcement learning with a misspecified linear function approximation. Our algorithm is parameter-free and does not require prior knowledge of misspecification level $\zeta$ or the suboptimality $\Delta$. Our algorithm is based on a novel certified estimator and provides the first constant regret guarantee for misspecified linear MDPs and (well-specified) linear MDPs. There are still some future works resulting from current algorithm. First, it is still an open question that whether the dependence on the planning horizon and dimension $d,H$ is optimal for instance-dependent regret bound, as in the well-specified case, the regret lower bound is $\Omega(d\sqrt{H^{3}K})$ (Zhou et al., 2021a), which has been recently attained by He et al. (2022a); Agarwal et al. (2022). Second, our work propose a new open question that if it possible to fill the gap between the positive result in our work and the negative result in Du et al. (2019). We believe it’s important in both theory and practice to understand how much misspecification level can be tolerated to efficiently learn the algorithm.
@@ Appendix A Additional Related Work
Recent years have witnessed a line of work focusing on RL with linear function approximation to tackle RL tasks in large state space. A widely studied MDP model is linear MDP (Jin et al., 2020), where both the transition kernel and the reward function are linear functions of a given feature mapping of the state-action pairs $\bm{\phi}(s,a)$. Several works have developed RL algorithms with polynomial sample complexity or sublinear regret bound in this setting. For example, LSVI-UCB (Jin et al., 2020) has an $\widetilde{\mathcal{O}}(\sqrt{d^{3}H^{4}K})$ regret bound, while randomized LSVI (Zanette et al., 2020a) has an $\widetilde{\mathcal{O}}(\sqrt{d^{4}H^{5}K})$ regret bound. He et al. (2022a) then improves this regret bound to a nearly minimax-optimal result $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ while Agarwal et al. (2022) provides a general function approximation extension given the above result. Linear mixture/kernel MDPs (Modi et al., 2020; Jia et al., 2020; Ayoub et al., 2020; Zhou et al., 2021b) have also emerged as another model that enables model-based RL with linear function approximation. In this setting, the transition kernel is a linear function of a feature mapping on the triplet of state, action, and next state $\bm{\phi}(s,a,s^{\prime})$. Nearly minimax optimal regrets can be achieved for both finite-horizon episodic MDPs (Ayoub et al., 2020; Zhou et al., 2021a) and infinite-horizon discounted MDPs (Zhou et al., 2021b) under this assumption.
@@ Appendix C Constant Regret Guarantees for Cert-LSVI-UCB
$\widecheck{V}^{k}_{h,l}(s)$ Pessimistic state value function in phase $l$. (see Algorithm 2)
@@
$\widecheck{V}^{k}_{h,l}(s)\leq\widehat{V}^{k}_{h}(s)\leq\widehat{V}^{k}_{h,l}(s),\quad|\widehat{V}^{k}_{h}(s)-\widehat{V}^{k}_{h,l}(s)|\leq 6\cdot 2^{-l}.$
@@ Appendix D Proof of Lemmas in Appendix C
According to the criteria for Line 11, we have $\widecheck{V}^{k}_{h,l}(s)\leq\widehat{V}^{k}_{h,l}(s)$ for any $l\in[l^{k}_{h}(s)-1]$. From the definition of $\widecheck{V}^{k}_{h,l}(s)$ and $\widehat{V}^{k}_{h,l}(s)$, they are monotonic in $l$ that $\widehat{V}^{k}_{h,l-1}(s)\leq\widehat{V}^{k}_{h,l}(s)$ and $\widehat{V}^{k}_{h,l}(s)\leq\widehat{V}^{k}_{h,l-1}(s)$ hold. Combining with $\widehat{V}^{k}_{h+1}(s)=\widehat{V}^{k}_{h,l^{k}_{h}(s)-1}$, we have
$\forall l\in[l_{h}^{k}(s)-1],\widecheck{V}^{k}_{h,l}(s)\leq\widehat{V}^{k}_{h}(s)\leq\widehat{V}^{k}_{h,l}(s)$ (D.1)
From the definition of $\widehat{V}^{k}_{h,l}(s)$ and $\widecheck{V}^{k}_{h,l}(s)$, we have
$0\leq\widehat{V}^{k}_{h,l}(s)-\widecheck{V}^{k}_{h,l}(s)\leq\big{(}\widehat{V}^{k}_{h,l}(s)-V^{k}_{h,l}(s)\big{)}+\big{(}V^{k}_{h,l}(s)-\widecheck{V}^{k}_{h,l}(s)\big{)}\leq 6\cdot 2^{-l}.$ (D.2)
@@
The following two lemmas demonstrate that, at stage $h$, both the optimistic state value function, $\widehat{V}^{k}_{h,l}(s)$, and the pessimistic state value function, $\widecheck{V}^{k}_{h,l}(s)$, exhibit a gap relative to the state value function determined by the Bellman operator, given as $V(s)=\max_{a\in\mathcal{A}}[\mathbb{B}_{h}\widehat{V}^{k}_{h+1}](s,a)$.
@@
$\max_{a\in\mathcal{A}}[\mathbb{B}_{h}\widehat{V}^{k}_{h+1}](s,a)-\max\big{\{}V^{k}_{h,l}(s)-3\cdot 2^{-l},\widecheck{V}^{k}_{h,l-1}(s)\big{\}}$ $\geq 2^{-l}-\chi\sqrt{L_{0}}\zeta,$
@@
$\max_{a\in\mathcal{A}}[\mathbb{B}_{h}\widehat{V}^{k}_{h+1}](s,a)-\widecheck{V}^{k}_{h,l}(s)$ $\geq 2^{-l}-\chi\sqrt{L_{0}}\zeta.$
@@
$\quad\min\big{\{}V^{k}_{h,l}(s)+3\cdot 2^{-l},\widehat{V}^{k}_{h,l-1}(s)\big{\}}-\max\big{\{}V^{k}_{h,l}(s)-3\cdot 2^{-l},\widecheck{V}^{k}_{h,l-1}(s)\big{\}}$
$=\big{(}\widehat{V}^{k}_{h,l}(s)-\max_{a\in\mathcal{A}}[\mathbb{B}_{h}\widehat{V}^{k}_{h+1}](s,a)\big{)}+\big{(}\max_{a\in\mathcal{A}}[\mathbb{B}_{h}\widehat{V}^{k}_{h+1}](s,a)-\widecheck{V}^{k}_{h,l}(s)\big{)}$
@@
whenever $\textrm{Regret}(K)>1$. Since $x\leq a+\sqrt{bx}$ implies $x\leq 2a+2b$, absorbing the case $\textrm{Regret}(K)\leq 1$ into the $\widetilde{\mathcal{O}}(\cdot)$ factor yields

### v2

@@ 6 Technical Challenges and Highlight of Proof Techniques
Our approach: A fine-grained concentration analysis
@@
We address this challenge by providing a fine-grained concentration analysis in connecting the gap with the regret. Notice that the regret $V^{*}_{h}(s_{h})-V^{\pi^{k}}_{h}(s_{h})$ in episode $k$ is the expectation of cumulative suboptimality gap $\operatorname*{\mathbb{E}}[\sum_{h=1}^{H}\Delta^{k}_{h}]$ taking over trajectory $\{s^{k}_{h}\}_{h=1}^{H}$. In addition, the variance of the random variable can be self-bounded according to
@@
Denote $\eta_{k}$ be the difference between $V^{*}_{h}(s_{h})-V^{\pi^{k}}_{h}(s_{h})$ and the actual $\sum_{h=1}^{H}\Delta^{k}_{h}$. Freedman inequality (Lemma H.5) implies that $\sum_{t=1}^{T}\eta^{t}\geq aC$ and $\sum_{t=1}^{T}\operatorname{{\rm Var}}[\eta^{t}]\leq vC$ happens at the same with a small probability for certain constant $a$ and $v$. Using a fine-grained union bound statement over $C$, we can reach the following statement indicates the cumulative regret can be upper bounded using the cumulative suboptimality gap:
@@
$\textstyle{{\sum_{k=1}^{K}}\big{(}V^{*}_{h}(s_{h})-V^{\pi^{k}}_{h}(s_{h})\big{)}\leq\widetilde{\mathcal{O}}\Big{(}{\sum_{k=1}^{K}}\sum_{h=1}^{H}}\Delta^{k}_{h}+H^{2}\Big{)}.$
Comparing with Lemma 6.1 in He et al. (2021a), Lemma 6.6 eliminates the $\log K$ dependence, which is achieved by the aforementioned fine-grained union bound. As a result, together with Lemma 6.4, we reach the desired statement that Cert-LSVI-UCB achieves constant regret bound when the misspecification is sufficiently small against the minimal suboptimality gap.
7 Conclusions and Limitations
In this work, we proposed a new algorithm, called certified estimator, for reinforcement learning with a misspecified linear function approximation. Our algorithm is parameter-free and does not require prior knowledge of misspecification level $\zeta$ or the suboptimality $\Delta$. Our algorithm is based on a novel certified estimator and provides the first constant regret guarantee for misspecified linear MDPs and (well-specified) linear MDPs.
Limitations.
Despite these advancements, several aspects of our algorithm and analysis warrant further investigation. One significant open question is whether the dependency on the planning horizon and dimension $d,H$ can achieve optimal instance-dependent regret bounds. For the gap-independent regret bounds, the regret lower bound is $\Omega(d\sqrt{H^{3}K})$ as shown by Zhou et al. (2021a), and this benchmark has recently been met by works such as He et al. (2022a); Agarwal et al. (2022). Additionally, our analysis assumes uniform misspecification across all actions. Investigating other types of misspecifications could lead to more sophisticated results, enhancing the algorithm’s robustness and applicability to diverse real-world scenarios. This exploration remains an important direction for future research.
Acknowledgments and Disclosure of Funding
We thank the anonymous reviewers for their helpful comments. This work was done while WZ was a PhD student at UCLA. WZ is partially supported by UCLA dissertation year fellowship and the research fund from UCLA-Amazon Science Hub. JH and QG are partially supported by the research fund from UCLA-Amazon Science Hub. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ Appendix A Additional Related Work
Recent years have witnessed a line of work focusing on RL with linear function approximation to tackle RL tasks in large state space. A widely studied MDP model is linear MDP (Jin et al., 2020), where both the transition kernel and the reward function are linear functions of a given feature mapping of the state-action pairs $\bm{\phi}(s,a)$. Several works have developed RL algorithms with polynomial sample complexity or sublinear regret bound in this setting. For example, LSVI-UCB (Jin et al., 2020) has an $\widetilde{\mathcal{O}}(\sqrt{d^{3}H^{4}K})$ regret bound, randomized LSVI (Zanette et al., 2020a) has an $\widetilde{\mathcal{O}}(\sqrt{d^{4}H^{5}K})$ regret bound and Ishfaq et al. (2021) achieved an $\widetilde{\mathcal{O}}(\sqrt{d^{3}H^{4}K})$. He et al. (2022a) then improves this regret bound to a nearly minimax-optimal result $\widetilde{\mathcal{O}}(d\sqrt{H^{3}K})$ while Agarwal et al. (2022) provides a general function approximation extension given the above result. Linear mixture/kernel MDPs (Modi et al., 2020; Jia et al., 2020; Ayoub et al., 2020; Zhou et al., 2021b) have also emerged as another model that enables model-based RL with linear function approximation. In this setting, the transition kernel is a linear function of a feature mapping on the triplet of state, action, and next state $\bm{\phi}(s,a,s^{\prime})$. Nearly minimax optimal regrets can be achieved for both finite-horizon episodic MDPs (Ayoub et al., 2020; Zhou et al., 2021a) and infinite-horizon discounted MDPs (Zhou et al., 2021b) under this assumption.
@@ Appendix C Constant Regret Guarantees for Cert-LSVI-UCB
[…]
@@
[…]
@@ Appendix D Proof of Lemmas in Appendix C
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
whenever $\textrm{Regret}(K)>1$. Taking $x=\textrm{Regret}(K)$, $a=\sum_{k=1}^{K}\sum_{h=1}^{H}\Delta^{k}_{h}+H^{2}\log(\textrm{Regret}(K)/\delta)$, and $b=2H^{2}\log(\textrm{Regret}(K)/\delta)$, inequality (1) yields
$\textrm{Regret}(K)$ $\leq 2\sum_{k=1}^{K}\sum_{h=1}^{H}\Delta^{k}_{h}+2H^{2}\log(\textrm{Regret}(K)/\delta)+4H^{2}\log(\textrm{Regret}(K)/\delta)$
$=2\sum_{k=1}^{K}\sum_{h=1}^{H}\Delta^{k}_{h}+6H^{2}\log(1/\delta)+6H^{2}\log(\textrm{Regret}(K))\quad$ (D.22)
According to the fact that $x\leq a\log x+b\Rightarrow x\leq 4a\log(2a)+2b$, letting $x=\textrm{Regret}(K),a=2\sum_{k=1}^{K}\sum_{h=1}^{H}\Delta^{k}_{h}+6H^{2}\log(1/\delta)$ and $b=6H^{2}$, (D.22) becomes
$\textrm{Regret}(K)\leq\left(8\sum_{k=1}^{K}\sum_{h=1}^{H}\Delta^{k}_{h}+24H^{2}\log(1/\delta)\right)\log\left(4\sum_{k=1}^{K}\sum_{h=1}^{H}\Delta^{k}_{h}+12H^{2}\log(1/\delta)\right)+12H^{2}.$
Hiding the logarithmic factors within the $\widetilde{O}$ notation yields